UseGauss-Jordan method to solve the following system of linear equations 2x – 3y + z = 0 5x + 4y + z = 10 2x – 2y – z = -1
Answer #1

22-3y + z = 0 5x+4y+z=10 2x-2y- Z = -1 From the expressed systens given above, it can be [mateix equation, as AX = В 2 - 3 102-310 -1 211 TO R3---3 R3 0 0 -431-43 s -3 Ilo -12 O Til is dark [A (13] = Rank(A) = 3.: The system consistent and since the